Find the orthagonal trijectries of the family of parabolas x=cy^2 and sketch their graph
1
Expert's answer
2019-09-30T11:04:08-0400

Differeniate "x=cy^2": "dx=2cydy" , so "y'=\frac{dy}{dx}=\frac{1}{2cy}"

Express "c" from the initial equation: "c=\frac{x}{y^2}". Then we have the differential equation of the family "x=cy^2" :"y'=\frac{1}{2cy}=\frac{1}{2y}\cdot\frac{y^2}{x}=\frac{y}{2x}=f(x,y)".

So the differential equation of the orthogonal family is "y'=-\frac{1}{f(x,y)}=-\frac{2x}{y}"

Solve this equation:

"\frac{dy}{dx}=-\frac{2x}{y}"

"ydy+2xdx=0"

"d\bigl(\frac{y^2}{2}+x^2\bigr)=0"

"\frac{y^2}{2}+x^2=C"

We obtain the family of ellipses "\frac{y^2}{2}+x^2=C=D^2" or "\frac{x^2}{D^2}+\frac{y^2}{2D^2}=1" "\frac{x^2}{D^2}+\frac{y^2}{2D^2}=1"

Answer: "\frac{x^2}{D^2}+\frac{y^2}{2D^2}=1"